OpenStudy (niahloves):

Kim has a small container and a large container. It takes 16 of the small container to fill the large container. Three small containers leave 1.95 gallons of space in the large container. What is the size of each of the two containers? A. The small container holds 2 gallons; the large container holds 15 gallons. B. The small container holds 0.15 gallons; the large container holds 2.4 gallons. C. The small container holds 2.4 gallons; the large container holds 150 gallons. D. The small container holds 15 gallons; the large container holds 24 gallons

OpenStudy (niahloves):

@GabeBae

OpenStudy (mathstudent55):

This is what the problem tells us: It takes 16 of the small container to fill the large container. 16 small = 1 large Three small containers leave 1.95 gallons of space in the large container. 3 small + 1.95 gal = 1 large Do you understand it so far?

OpenStudy (niahloves):

Kinda @mathstudent55

OpenStudy (mathstudent55):

The first statement is very simple. 16 small containers = 1 large container. I don't think you have any questions on it, right?

OpenStudy (babylove18):

its b sorry i did it on the other post

OpenStudy (babylove18):

2.4-1.95=.45/.15=3

OpenStudy (babylove18):

@niahloves

OpenStudy (mathstudent55):

The second statement is that if you pour 3 small containers into the large one, you are still short, by 1.95 gal, of filing the large container That means 3 small containers + 1.95 gallons = the large container

OpenStudy (mathstudent55):

Once again we have: 16 small = 1 large 3 small + 1.95 gal = 1 large But 16 small = 3 small + 13 small, so we substitute: 3 small + 13 small = 1 large 3 small + 1.95 gal = 1 large Now when we compare the last two equations above, we see that 13 small containers = 1.95 gal. small container = (1.95 gal)/13 = 0.15 gal The small container measure 0.15 gal. The large container is 16 times larger, so large container = 16 * 0.15 gal = 2.4 gal Answer: small container 0.15 gal; large container 2.4 gal